How many solutions does the system 3x-y=-6 & 3x+y=6 have?
Answer by jim_thompson5910(35256)

Add up the equations to get

3x-y=-6
3x+y=6
-------
6x+0y = 0

That leads to 6x = 0 which means x = 0. You then use x = 0 to find y. Plug it into any equation with x & y in it. You'll find that y = 6.

So there is exactly one solution and it is the ordered pair (0,6)